# Binarizing Physics-Inspired GNNs for Combinatorial Optimization

### v1

@@ 4 Experiments
The experiments were aligned with the original methodology of PI-GNNs [47] by choosing equivalent CO problems, data collection approach, and overall research setup. The purpose of the experiments is to address the problematic PI-GNN training performance for denser CO settings (Sec. 3.1) with the methods introduced in Sec. 3.2 and 3.3.
@@ Methods
Particularly, apart from the original PI-GNN approach (baseline), we propose a portfolio of the following $7$ methods based on the concepts of (i) soft (annealing) and (ii) hard binarization (Sec. 3.3), and (iii) fuzzy relaxation (Sec. 3.2):
@@
For a direct comparison with Schuetz et al. [47], we make use of the same GNN ansatz, with each training experiment being repeated over $N=5$ random initializations. The GNN model is iteratively trained on individual graph instances against a custom loss function that encodes a relaxed version of the optimization problem, as specified by the Hamiltonian $H_{Q}$ (Fig. 5). The optimization hyperparameters are maintained from Schuetz et al. [47], using the Adam optimizer [29] and early stopping in case of no improvement.
@@
Aiming at improving the performance of PI-GNNs over graphs of increasing connectivity, we vary the density of the graphs $\mathcal{G}$, controlled by the regularity parameter $d\in\{3,5,10,20,30,40,50\}$. Otherwise, the data collection routine remains identical to the one introduced by Schuetz et al. [47], generating $20$ random graphs for each graph type (combination of problem and $d$-regularity).
@@ 4.1 Results
The portfolio of the proposed methods, using different discretization and fuzzification techniques, outperformed the baseline QUBO on all datasets, measured with both best-of-N ($BoN$) and average ($Avg$) metrics. See Tab. 1 and Tab. 2 for $BoN$ in the MaxCut and MIS domains. Specifically, the soft discretization techniques based on temperature annealing (temp-lin, temp-log, temp-exp) slightly outperformed the baseline on sparse graphs, while the binarization and fuzzy-logic adjustments clearly dominated all the denser setups. Similar results then prevail when upscaling the graphs.
@@
Many of the proposed architectures also exhibited improved training dynamics (Sec. 3.1). Specifically, the pre-activations of binarized sigmoids (bin-ste and bin-sig) remained appropriately in the proximity of zero, and the Łukasiewicz conjunction also converged more smoothly in most cases compared to the baseline. The pre- and post-activation values for bin-ste and fuzzy-luk are depicted in Fig. 6.
For a complete overview of training hyperparameters, scaling results with varying graph sizes $n\in\{100,1000,10000\}$, and evaluation with the $Avg$ metric, see supplementary materials. The source code is available at https://github.com/martin-krutsky/gnn-comb-opt.
@@ Appendix A MIS Results
As the main text, due to space constraints, concentrated on presenting the results on the maximum cut problem, we report the results on the maximum independent set in Tab. 3 and 4. Similarly to MaxCut, the proposed discretization and fuzzy-logical adjustments outperform the baseline.
@@ Appendix B Scaling the Experiments
To compare the behavior of the proposed methods over graphs of increasing size, similarly to Schuetz et al. [48], we present results with respect to the best-of-N metric on graphs with $100$, $1000$, and $10000$ nodes in Tab. 5, 6 and 7. Due to the size of our experimental grid and resources needed, the experiments on the largest graphs of size $N=10000$ were restricted to the following settings for the MaxCut problem: we tested all of our $8$ candidate architectures on single random graphs of regularities $d\in\{3,5,10,20,30,40,50\}$ with $5$ random initializations. The results show that the discretization methods remain a promising improvement with the scaling of the size of the input.
@@ Appendix C Results by Average Performance
The main text presents results based on the best-of-N (BoN) aggregation, as it is the preferred one with respect to the parallelization of the training [47]. For completeness, see the results for $N\in\{100,1000\}$ based on averaging (Avg) metric in Tab. 8, 9, 10 and 11. The averaging metric requires adjustments for non-feasible solutions (which are possible only for MIS). While in the BoN metrics we could disregard such results as long as at least one of the seeds for each setting resulted in a feasible solution, for Avg, infeasible results should logically affect the final arithmetic mean. We nullified such results, with the intuition that the bitstring of zeros represents the trivial result of empty maximal sets and cuts.
@@ Appendix D Comparison to Solvers
Similarly to previous work, the results on smaller instances have been compared with solvers. Notably, for MaxCut, the Goemans-Williamson algorithm [21] has been used. The solver can, however, handle only graphs with a couple of hundred nodes in a reasonable time and with reasonable memory demands [47]. On data with $N=100$ nodes we have observed: (i) comparable results on the low-regularity graphs, retrieving $0.97$ to $0.98$ of the solver’s prediction value with our improved PI-GNN approach; (ii) lower results on the high-regularity graphs, retrieving $0.71$ to $0.72$ of the solver’s prediction value with our improved PI-GNN approach. Thus, while our approach surpasses the original PI-GNNs proposal [47] and, unlike classical solvers, remains efficient on larger instances, improving PI-GNNs’ performance on small graphs, over classical solvers, remains open for further exploration.
@@ Appendix E Regularization Ablation
Concretely, two regularizations were considered: the L1 penalty: $P(x)=\sum_{i=1}^{N}|x_{i}|$, and binary entropy: $P(x)=\sum_{i=1}^{N}[-x_{i}\cdot log_{2}(x_{i})-(1-x_{i})\cdot log_{2}(1-x_{i})]$. The former is more directed towards sparsity, as it encourages values close to zero. The latter is suited directly for binary discretization, being symmetric for 00 and $1$.
@@ Appendix F Hyperparameters
Training and GNN hyperparameters are disclosed in Tab. 12. The settings, following the methodology of Schuetz et al. [47], were shared over all the experiments.
@@
Number of Layers Total number of graph convolutional layers $2$
@@ Appendix G Heatmaps of Node Assignment Dynamics
Further heatmaps depicting the training dynamics for $5$-regular MaxCut and MIS graphs are in Fig. 7 and 8. Similarly to graphs in the main text, we can see the phase transition in the upper right corners and improved training dynamics in the second and third rows.
Figure 7: Distributions of GNN’s pre-activation ($a^{pre}$) and post-activation ($a^{post}$) values on $5$-regular MaxCut graphs: (a) baseline pre-activations, (b) baseline post-activations, (c) sign-ste pre-activations, (d) sign-ste post-activations, (e) fuzzy-luk pre-activations, (f) fuzzy-luk post-activations.
Figure 8: Distributions of GNN’s pre-activation ($a^{pre}$) and post-activation ($a^{post}$) values on $5$-regular MIS graphs: (a) baseline pre-activations, (b) baseline post-activations, (c) sign-ste pre-activations, (d) sign-ste post-activations, (e) fuzzy-luk pre-activations, (f) fuzzy-luk post-activations.

### v2

@@ 4 Experiments
The experiments were aligned with the original methodology of PI-GNNs [49] by choosing equivalent CO problems, data collection approach, and overall research setup. The purpose of the experiments is to address the problematic PI-GNN training performance for denser CO settings (Sec.˜3.1) with the methods introduced in Sec.˜3.2 and 3.3.
@@ Methods
Particularly, apart from the original PI-GNN approach (baseline), we propose a portfolio of the following 777 methods based on the concepts of (i) soft (annealing) and (ii) hard binarization (Sec.˜3.3), and (iii) fuzzy relaxation (Sec.˜3.2):
@@
For a direct comparison with Schuetz et al. [49], we make use of the same GNN ansatz, with each training experiment being repeated over $N=5$ random initializations. The GNN model is iteratively trained on individual graph instances against a custom loss function that encodes a relaxed version of the optimization problem, as specified by the Hamiltonian $H_{Q}$ (Fig.˜5). The optimization hyperparameters are maintained from Schuetz et al. [49], using the Adam optimizer [30] and early stopping in case of no improvement.
@@
Aiming at improving the performance of PI-GNNs over graphs of increasing connectivity, we vary the density of the graphs $\mathcal{G}$, controlled by the regularity parameter $d\in\{3,5,10,20,30,40,50\}$. Otherwise, the data collection routine remains identical to the one introduced by Schuetz et al. [49], generating 202020 random graphs for each graph type (combination of problem and $d$-regularity).
@@ 4.1 Results
The portfolio of the proposed methods, using different discretization and fuzzification techniques, outperformed the baseline QUBO on all datasets, measured with both best-of-N ($BoN$) and average ($Avg$) metrics. See Tab.˜1 and Tab.˜2 for $BoN$ in the MaxCut and MIS domains. Specifically, the soft discretization techniques based on temperature annealing (temp-lin, temp-log, temp-exp) slightly outperformed the baseline on sparse graphs, while the binarization and fuzzy-logic adjustments clearly dominated all the denser setups. Similar results then prevail when upscaling the graphs.
@@
Many of the proposed architectures also exhibited improved training dynamics (Sec.˜3.1). Specifically, the pre-activations of binarized sigmoids (bin-ste and bin-sig) remained appropriately in the proximity of zero, and the Łukasiewicz conjunction also converged more smoothly in most cases compared to the baseline. The pre- and post-activation values for bin-ste and fuzzy-luk are depicted in Fig.˜6.
For a complete overview of training hyperparameters, scaling results with varying graph sizes $n\in\{100,1000,10000\}$, and evaluation with the $Avg$ metric, see App.˜A, B, C and F. The source code is available at https://github.com/martin-krutsky/gnn-comb-opt.
@@ Appendix A MIS Results
As the main text, due to space constraints, concentrated on presenting the results on the maximum cut problem, we report the results on the maximum independent set in Tab.˜3 and 4. Similarly to MaxCut, the proposed discretization and fuzzy-logical adjustments outperform the baseline.
@@ Appendix B Scaling the Experiments
To compare the behavior of the proposed methods over graphs of increasing size, similarly to Schuetz et al. [50], we present results with respect to the best-of-N metric on graphs with 100100100, 100010001000, and 100001000010000 nodes in Tab.˜5, 6 and 7. Due to the size of our experimental grid and resources needed, the experiments on the largest graphs of size $N=10000$ were restricted to the following settings for the MaxCut problem: we tested all of our 888 candidate architectures on single random graphs of regularities $d\in\{3,5,10,20,30,40,50\}$ with 555 random initializations. The results show that the discretization methods remain a promising improvement with the scaling of the size of the input.
@@ Appendix C Results by Average Performance
The main text presents results based on the best-of-N (BoN) aggregation, as it is the preferred one with respect to the parallelization of the training [49]. For completeness, see the results for $N\in\{100,1000\}$ based on averaging (Avg) metric in Tab.˜8, 9, 10 and 11. The averaging metric requires adjustments for non-feasible solutions (which are possible only for MIS). While in the BoN metrics we could disregard such results as long as at least one of the seeds for each setting resulted in a feasible solution, for Avg, infeasible results should logically affect the final arithmetic mean. We nullified such results, with the intuition that the bitstring of zeros represents the trivial result of empty maximal sets and cuts.
@@ Appendix D Comparison to Solvers
Similarly to previous work, the results on smaller instances have been compared with solvers. Notably, for MaxCut, the Goemans-Williamson algorithm [21] has been used. The solver can, however, handle only graphs with a couple of hundred nodes in a reasonable time and with reasonable memory demands [49]. On data with $N=100$ nodes we have observed: (i) comparable results on the low-regularity graphs, retrieving 0.970.970.97 to 0.980.980.98 of the solver’s prediction value with our improved PI-GNN approach; (ii) lower results on the high-regularity graphs, retrieving 0.710.710.71 to 0.720.720.72 of the solver’s prediction value with our improved PI-GNN approach. Thus, while our approach surpasses the original PI-GNNs proposal [49] and, unlike classical solvers, remains efficient on larger instances, improving PI-GNNs’ performance on small graphs, over classical solvers, remains open for further exploration.
@@ Appendix E Regularization Ablation
Concretely, two regularizations were considered: the L1 penalty: $P(x)=\sum_{i=1}^{N}|x_{i}|$, and binary entropy: $P(x)=\sum_{i=1}^{N}[-x_{i}\cdot log_{2}(x_{i})-(1-x_{i})\cdot log_{2}(1-x_{i})]$. The former is more directed towards sparsity, as it encourages values close to zero. The latter is suited directly for binary discretization, being symmetric for 0 and 111.
@@ Appendix F Hyperparameters
Training and GNN hyperparameters are disclosed in Tab.˜12. The settings, following the methodology of Schuetz et al. [49], were shared over all the experiments.
@@
Number of Layers Total number of graph convolutional layers 222
@@ Appendix G Heatmaps of Node Assignment Dynamics
Further heatmaps depicting the training dynamics for 555-regular MaxCut and MIS graphs are in Fig.˜7 and 8. Similarly to graphs in the main text, we can see the phase transition in the upper right corners and improved training dynamics in the second and third rows.
Figure 7: Distributions of GNN’s pre-activation ($a^{pre}$) and post-activation ($a^{post}$) values on 555-regular MaxCut graphs: (a) baseline pre-activations, (b) baseline post-activations, (c) sign-ste pre-activations, (d) sign-ste post-activations, (e) fuzzy-luk pre-activations, (f) fuzzy-luk post-activations.
Figure 8: Distributions of GNN’s pre-activation ($a^{pre}$) and post-activation ($a^{post}$) values on 555-regular MIS graphs: (a) baseline pre-activations, (b) baseline post-activations, (c) sign-ste pre-activations, (d) sign-ste post-activations, (e) fuzzy-luk pre-activations, (f) fuzzy-luk post-activations.